General Announcement: - Construction noises heard at XHQ 2nd floor comes from HXS works at the experimental Hall. - Tobias Clever: a power glitch was seen on Sunday around 17:00 and some pumps supplying cooling water to XHQ level 0 were affected; also ventilation at the buildings was affected. TS OCD could restore the pumping system before temperatures have rising. CTRL: - Proceeding well with Karabo3 tests at MID. DA: - Nothing to report. EEE: - Continue with maintenance at SA2 PLC systems. - SA3 undulators will be also upgrades. VAC: - Nothing to report. XPD: - Nothing to report; everything ongoing as planned; all normal with XGM pressure. DET: - Nothing to report; AGIPD 4M: to be discussed offline. ITDM: - Nothing to report. XRO: - Nothing to report. LAS: - Not present. UNSYS: - Everything is fine; installation in SA2 is ongoing. SPB: - Nothing to report. FXE: - Not present. MID: - All fine at MID. HED: - Nothing to report. SCS: - Not present - nothing to report. SQS: - Nothing to report. SXP: - Nothing to report.